### v1

# FiDeLiS: Faithful Reasoning in Large Language Model for Knowledge Graph Question Answering

Abstract
While large language models (LLMs) have achieved significant success in various applications, they often struggle with hallucinations, especially in scenarios that require deep and responsible reasoning. These issues could be partially mitigated by integrating external knowledge graphs (KG) in LLM reasoning. However, the method of their incorporation is still largely unexplored. In this paper, we propose a retrieval-exploration interactive method, FiDeLiS to handle intermediate steps of reasoning grounded by KGs. Specifically, we propose Path-RAG module for retrieving useful intermediate knowledge from KG for LLM reasoning. We incorporate the logic and common-sense reasoning of LLMs and the topological connectivity of KGs into the knowledge retrieval process, which provides more accurate retrieving performance. Furthermore, we propose to leverage deductive reasoning capabilities of LLMs as a better criterion to automatically guide the reasoning process in a stepwise and generalizable manner. Deductive verification serves as a precise indicator for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation. Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks. The code of this paper will be released later.
1 Introduction
The emergence and application of large language models (LLMs) [1, 31, 27] have attracted widespread attention from researchers and the general public. It demonstrates remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning [47, 49]. However, the reasoning of LLMs is not invariably reliable and may conflict with factual reality [32], which limits the application in areas requiring high reliability, such as healthcare [14] and science [41].
The knowledge graph (KG) stores high-quality domain-specific knowledge in structured triplets, such as Wikidata [43], YAGO [36], and NELL [2]. KGs offer structured, explicit, and updated factual knowledge, which provides a faithful knowledge source for reasoning. Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretive process transparent. Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning. Therefore, several attempts have been conducted to integrate KGs with language models [32, 29, 17, 38]. Among them, the knowledge graph question answering (KGQA) is a critical task to incorporate the knowledge of KG into reasoning models [12, 44, 55]. Concretely, KGQA faces the following main challenges:
(I) How to retrieve specific knowledge from KGs to help reasoning precisely? Existing solutions include direct retrieval [37, 21] and semantic parsing [39, 10]. Direct retrieval involves taking the questions as a query and the knowledge triplets in the KG as candidates, using either sparse or dense retrieval techniques to identify several candidates most relevant to the query. Semantic parsing, on the other hand, transforms the question into an executable structured query statement (e.g., SPARQL) and executes the query in KGs. However, in KGs, some triplet by itself lacks rich semantic information due to the difference of KG schema, such as some entities are represented as machine identifiers (MIDs), and direct retrieval make it difficult to understand how relevant it really is to the user query. This becomes especially tricky in multi-hop question answering, where the intermediate triplets seem not directly related to the query can actually be crucial links to finding the right answer. Furthermore, semantic parsing faces the issues of non-executable or incorrectly executed generated queries [55].
(II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs? Existing solutions include (1) finetune LMs on KG structure to generate relation path grounded by KGs. These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results [29, 55]; (2) leverage the reasoning model to iteratively retrieve and reason on the subgraphs from KG [38, 9, 21] (i.e., at each step, the LLMs retrieve neighbors of the current nodes and then decide to answer the question or continue the next search step.) However, for the first approach, the reasoning steps generated by language models offer few guarantees about feasibility (i.e., the reasoning step might not exist in the KG), especially when multiple consecutive steps are combined into a reasoning path. We conduct the error analysis regarding this issue in Section 4. The latter solution face the challenge regarding determining the optimal stopping point for the exploration process. For example, ToG [38] prompts the LLMs to evaluate the adequacy of current reasoning paths for answer generation. However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain. This lack of clarity can easily result in either premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
To address these two challenges, we propose a retrieval-exploration interactive method, FiDeLiS. It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process. Specifically, we propose a Path-RAG module for retrieving useful intermediate knowledge from KGs. Compared to vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked. We also integrate the topological connectivity of KGs to enhance system accuracy by considering entity relationships in the knowledge retrieval process. Furthermore, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion for automatically directing the reasoning process step by step. Deductive reasoning serves as a clear indicator for ceasing further reasoning when necessary, which helps prevent errors in the reasoning chain and reduce unnecessary computation. Our framework enables us to employ beam search, a technique tailored for intermediate reasoning steps within a controllable searching space. We can guide the searching process by ensuring the logical soundness of each reasoning step to prevent errors throughout the reasoning process. Overall, our main contributions include:
  • We propose a retrieval-exploration interactive method that is specifically tailored to handle intermediate steps of reasoning grounded by KGs.
  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and unnecessary computation.
  • We integrate the reasoning process and topological connectivity of KGs into the knowledge retrieval process, which aids in recalling useful intermediate knowledge for reasoning and maintains efficiency as KGs scale up.
  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality outperforms the existing strong baselines in three benchmarks.
2 Method
Overview. FiDeLiS is designed as a retrieval-exploration interactive method as shown in Figure 1. It is designed to iteratively retrieve and reason on the subgraphs from KG, simulating a dynamic searching process through KG. FiDeLiS incorporates two modules Reasoning Path Retrieval-Augmented Generation (Path-RAG) and Deductive-Verification Guided Beam Search (DVBS), where Path-RAG is utilized to retrieve a highly pertinent set of nodes/relations from KGs to the user query and determine the reasoning step based on the relevance score discussed in Section 2.1. On the other hand, DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension. To prevent either premature stopping or excessive continuation of reasoning path extension, DVBS enhances the clarity of termination signals using deductive reasoning discussed in Section 2.2.
2.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Path-RAG comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Initialization. We initiate the Path-RAG by generating node and edge embeddings using a pre-trained Language Model (LM). These embeddings are then stored in a nearest neighbor data structure. To elaborate, consider $e_{i}\in\mathcal{E},r_{i}\in\mathcal{R}$, where $\mathcal{E}$ refers to the entities and $\mathcal{R}$ represents the relations in the relational graph. We leverage the pre-trained LM (e.g., SentenceBert [33] and E5 [46]) to $e_{i}$ and $r_{i}$, yielding the representation $z(e^{i})$ and $z(r^{i})$:
$z(e^{i})=\operatorname{LM}\left(e_{i}\right)\in\mathbb{R}^{d},\\ z(r^{i})=\operatorname{LM}\left(r_{i}\right)\in\mathbb{R}^{d}$ (1)
where $d$ denotes the dimension of the output vector.
Retrieval. To retrieve the entities and relations that are relevant to the user query, the embeddings of them are then utilized to populate a nearest neighbor index. Specifically, given a query $q$, we leverage the LM to generate an exhaustive list of keywords or relation names $k_{i}\in\mathcal{K}$ that would be used to find the reasoning path from the knowledge graph to answer the query (as referred to prompt list in Section B). The generated list of keywords aims to maximize the coverage of potential reasoning steps and to enhance the subsequent decision-making process of LLMs by concentrating on a smaller set of relevant reasoning step candidates.
$\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$ (2)
Then each $k_{i}^{m}\in\mathcal{K}$ is concatenated into a string $k_{i}$ and embedded into the same latent space by the LM to obtain $z(k)\in\mathbb{R}^{d}$, the retrieval process involves identifying the node or sequence of nodes that best aligns with the semantic content of the keywords as follow:
$\mathcal{E}_{m}=\operatorname{argtopm}_{i\in\mathcal{E}}\cos\left(z(k),z(e)\right),\mathcal{R}_{m}=\operatorname{argtopm}_{i\in\mathcal{R}}\cos\left(z(k),z(r)\right)$ (3)
where $z(e)$ and $z(r)$ are the embeddings of nodes $e$ and edges $r$, respectively. We use the cosine similarity function, $\cos(\cdot,\cdot)$, to measure the similarity between the keyword representation and the node/edge embeddings. The $\operatorname{argtopm}$ operation retrieves the top-m elements based on this similarity, providing a set of nodes $e_{k}$ and edges $r_{k}$ considered most relevant to the query. We denote the score of each entity and each relation as $S_{e}^{i}=\cos\left(z(k),z(e)\right)$ and $S_{r}^{i}=\cos\left(z(k),z(r)\right)$.
Reasoning Step Candidates Construction. To construct reasoning steps candidates based on the finite set of nodes $\mathcal{E}_{m}$ and edges $\mathcal{R}_{m}$ relevant to the query and the relevance score $S_{e}^{i}$ and $S_{r}^{i}$, we denote the reasoning step can be scored and retrieved based on the following criteria as follows. The goal of this scoring function is to leverage the triplets to represent the connectivity of the KGs. It provides some hints on whether the current operation/decision would lead to a malicious operation several steps later by incorporating the next-hop neighbor’s information.
$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)$ (4)
It should be noted that the entity and relation mentioned in the equation should be logically connected in the corresponding KG. $S(p)$ denotes the score of the reasoning step; $S_{r}^{i}$, $S_{e}^{i}$ denote the relevance score of the next step relation and the corresponding neighbor; $S_{r}^{j}$, $S_{e}^{j}$ denote the relevance score of one further hop of next step. Parameter $\alpha$ is used to balance between immediate accuracy and long-term prospects within the graph. A higher $\alpha$ value encourages the model to consider less immediate but potentially vital outcomes, While a lower $\alpha$ makes it focus primarily on the immediate next steps without considerable regard for the long-term impacts. The detailed description can be found in Appendix Algorithm 1.
Figure 1: Overview of FiDeLiS. The goal of FiDeLiS is to extend the multiple-step reasoning paths grounded by the external knowledge graph. We instantiate the $T$ step reasoning path extension as shown in the figure. First, Path-RAG is used to retrieve a highly pertinent set of nodes/relations from KGs to the user query and construct the reasoning step candidates. Then, the constructed reasoning steps candidates are forwarded to DVBS. DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates for finding the most promising reasoning step. Specifically, at timestep $T$, DVBS leverage the LLMs to choose the top-$k$ (here we set the $k=2$ for better clarification) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on the deductive verification.
2.2 DVBS: Deductive-Verification Guided Beam Search
DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) for finding the most promising reasoning paths. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path extension. It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Plan-and-Solve. Inspired by the recent works regarding planning capabilities of LLMs [35, 56], this step aims to provide more hints for subsequent LLM decision-making. Specifically, we prompt LLM to generate the planning steps for searching the most promising reasoning paths initially denoted as $p$. The detailed prompts can be found in Section B.
Beam Search. In multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $R=\left[s^{1},s^{2},\ldots,s^{T}\right]=s^{1:T}$ as shown in Figure 1. However, different from the typical text decoding process where each step consists of a single token, we follow the [50] and view a sequence of pair $(r,e)$ as a single step, where $r$ is the relation and $e$ is the corresponding entity. At each timestep, we leverage the LLMs to choose the top-k reasoning steps and decide whether to continue the next search step or cease the reasoning path. We present the beam search at each step is modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{K}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\right\}$ (5)
where $\mathcal{H}_{t-1}$ denote the reasoning path at the previous time step $t-1$; $s^{t}$ refers to the current reasoning step; $\mathcal{S}$ refers to all possible next steps $s^{t}$ candidates produced from the Path-RAG; $\text{LM}(s^{t}|q,s^{1:t-1},p)$ refers to the language models prediction for the next step $s^{t}$ given the current sequence $s^{1:t-1}$, the query $q$, and the planning context $p$; $\oplus$ refers to appending $s^{t}$ to the current sequence $h$; $\text{Top}_{K}$ refers to selecting the top $K$ (denoted as beam width in the experiment setting) sets of reasoning steps.
Deductive Verification. To mitigate the reasoning path extension terminates at the wrong point, we propose to leverage deductive reasoning capabilities of LLMs [26, 18] as a better criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 5. Specially, we first leverage LLMs to convert the user query to a declarative statement (as shown in Appendix G.2) and then use LLMs to judge whether the declarative statement of query $x^{\prime}$ can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$. We utilize the same backend LLM with prompt referred in Section B and define the deductive-verification criteria as follows:
$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1&\text{if declarative statement ofquery }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0&\text{otherwise.}\end{cases}$ (6)
The overall DVBS can be modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},p):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (7)
This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning to make sure it follows logically and matches the original query. It acts as a clear sign to stop further reasoning when necessary, which helps avoid mistakes in the reasoning process and reduces unnecessary computations.
3 Experiments
We aim to answer the following research questions during this section: RQ1: How does FiDeLiS compare with existing baselines in terms of accuracy and efficiency on KGQA tasks? RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS? RQ3: How robustness of FiDeLiS perform under varying conditions?
3.1 Experiment Settings
Datasets & Metrics. We adopt three benchmark KGQA datasets: WebQuestionSP (WebQSP) [54], Complex WebQuestions (CWQ) [40] and CR-LT-KGQA [11] in this work. We follow previous work [29] to use the same training and testing splits for fair comparison over WebQSP and CWQ. Both WebQSP and CWQ can be reasoned using Freebase KGs [^1] https://github.com/microsoft/FastRDFStore. To address the bias in WebQSP and CWQ, which predominantly feature popular entities and there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, we tested our methods on CR-LT-KGQA (further discussed in Section 5). We use the complete dataset from CR-LT-KGQA in our experiments, as it comprises only 200 samples. Each of these can be reasoned based on the Wikidata KG [^2] https://www.wikidata.org/wiki/Wikidata:Main_Page. The statistics of the datasets are given in Table 8 and Table 9.
To streamline the KGs, we utilize a subgraph of Freebase by extracting all triples that fall within the maximum reasoning hops from the question entities in WebQSP and CWQ [29]. Similarly, we construct the corresponding sub-graphs of Wikidata for CR-LT-KGQA as well. We assess the performance of the method by analyzing the F1 and Hits@1 metrics for CWQ and WebQSP, and by evaluating the accuracy for CR-LT-KGQA.
Baselines. (1) RoG [29]: embed structure knowledge graph from KGs into the underlying neural networks during the pretraining and fine-tuning process to generate the reasoning path and explanation. (2) ToG [38]: ask LLM to iteratively explore multiple possible reasoning paths on KGs until the LLM determines that the question can be answered based on the current reasoning paths. Our method FiDeLiS follows this paradigm. (3) NSM [12] utilizes the sequential model to mimic the multi-hop reasoning process. (4) KD-CoT [44] retrieves relevant knowledge from KGs to generate faithful reasoning paths for LLMs. (5) DeCAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance on KGQA tasks. The details of each baseline are further discussed in Section G.1.
Implementation Details. We set the default beam width as $4$ and depth as $4$ without specific annotation. We set the $\alpha$ in Eq 4 as $0.3$ to ensure reproducibility.
Backboned LLMs. We assess our approach on closed- and open-source LLMs. For closed-source LLMs, we choose GPT-4-turbo, GPT-3.5-turbo [^3] We use the recent gpt-4-turbo model that is released in 2024-04-09 from https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 and gpt-3.5-turbo model notated as “gpt-3.5-turbo-0125” from https://platform.openai.com/docs/models/gpt-3-5-turbo. to report and compare the results on all datasets. We use Llama-2-13B [42] and Mistral-7B [19] as our open-source LLMs to conduct cost–performance analysis on different datasets. The experiment results of open-source models can be found in Table 10.
3.2 RQ1: KGQA Performance Comparison
Comparison between baseline methods. Table 1 clearly showcases the comparative effectiveness of FiDeLiS, against different baselines. It shows that FiDeLiS outperforms all baselines integrated with gpt-4-turbo, even for strong baselines being trained, such as DeCAF [55] and RoG [29]. Across all prompting-based methods, gpt-4-turbo achieves higher performance compared to gpt-3.5-turbo, especially on the CWQ dataset. It indicates that gpt-4-turbo has a better understanding and processing capability for complex queries. The CR-LT dataset seems more demanding than WebQSP and CWQ, as implied by the consistently lower accuracy scores. However, FiDeLiS shows consistent enhancements compared to other baselines, demonstrating its capability to handle long-tail entities by referring to the knowledge from KGs and processing more complex queries effectively.
Table 1: Comparison of FiDeLiS with baseline methods and different backbone LLMs. We replicate the outcomes of ToG and RoG, and retrieve other baseline results directly from the original paper. We utilize 5 demonstrations as our default setting for FiDeLiS, ToG, Few-shot, and CoT. The experiment results of open-source models can be found in Table 10.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Prompting - LLM Only gpt-3.5-turbo Zero-shot 54.37 52.31 34.87 28.32 32.74
Few-shot 56.33 53.12 38.52 33.87 36.61
CoT 57.42 54.72 43.21 35.85 37.42
Prompting - LLM Only gpt-4-turbo Zero-shot 62.32 59.71 42.71 37.93 37.74
Few-shot 68.65 62.71 51.52 43.70 43.61
CoT 72.11 65.37 53.51 44.76 45.42
Finetuning - LLM + KG NSM [12] 74.31 - 53.92 - -
CBR-KBQA [6] - - 67.14 - -
DeCAF [55] 82.1 - 70.42 - -
KD-CoT [44] 73.7 50.2 50.5 - -
RoG [29] 83.15 69.81 61.39 56.17 60.32
Prompting - LLM + KG gpt-3.5-turbo ToG [38] 75.13 72.32 57.59 56.96 62.48
FiDeLiS 79.32 76.78 63.12 61.78 67.34
Prompting - LLM + KG gpt-4-turbo ToG [38] 81.84 75.97 68.51 60.20 67.24
FiDeLiS 84.39 78.32 71.47 64.32 72.12
3.3 RQ2: Ablation Study of FiDeLiS.
Table 2 provides a detailed view of how removing specific components from FiDeLiS impacts its performance. The results emphasize that some features, such as Path-RAG and beam search in DVBS, are particularly significant, especially for tackling complex reasoning tasks on datasets like CWQ. When Path-RAG and beam-search within DVBS are excluded, we observe substantial declines in performance. This underlines their importance in our model’s ability to handle intricate reasoning effectively. Furthermore, the integration of the reasoning and planning steps within DVBS plays a critical role in maintaining robust performance across varied datasets. One aspect highlighted by our study is the role of deductive verification during the beam search phase in DVBS. This feature enhances the model’s ability to delineate and reinforce decision boundaries, thus safeguarding against reasoning errors.
Table 2: Ablation Studies of FiDeLiS, we test the ablation study using model gpt-3.5-turbo-0125. DVBS refers to “Deductive-verification guided Beam Search’. “$\Delta$” refers to the performance gap between each component and the entire method. The top-3 performance gaps on each dataset are highlighted in green, with darker shades indicating significant performance gap.
- Components WebQSP CWQ CR-LT
Hits@1 (%) $\Delta$ Hits@1 (%) $\Delta$ Acc (%) $\Delta$
- FiDeLiS 79.32 0.00 63.12 0.00 67.34 0.00
w/o Path-RAG using vanilla retriever 72.35 6.97 57.11 6.01 59.78 7.56
using ToG 75.11 4.21 59.47 3.65 63.47 4.07
w/o DVBS w/o last step reasoning 75.68 3.64 59.45 3.67 63.72 3.62
w/o planning 76.23 3.09 60.14 2.98 64.13 3.21
w/o beam-search 60.35 18.97 49.78 13.34 61.87 5.47
w/o deductive-verifier 74.13 5.19 57.23 5.89 63.89 3.45
3.4 RQ3: Robustness Analysis
Comparison using different embedding backbones. Table 3 demonstrates the performance of the Path-RAG using various embedding models as backbones. It shows that the integration of the Openai-Embedding Model significantly enhances the performance across all datasets, indicating its superior capability in understanding and processing natural language queries, which leads to more effective retrieval. For instance, the Openai-Embedding Model increases the Hits@1 by 13.04% in the WebQSP dataset and 14.73% in the CWQ dataset compared to the base BM25 [34] model. Furthermore, this improvement is consistent and highlights the feasibility and potential of pairing more powerful language models with structured reasoning frameworks to address complex information retrieval and question answering challenges.
Table 3: Performance FiDeLiS using different Path-RAG embedding backbone models. ‘Openai-Embedding Model’ refers to text-embedding-3-small, the recent embedding model released from OpenAI.
Methods Backbones WebQSP CWQ CR-LT
Hits@1 (%) Hits@1 (%) Acc (%)
Vanilla Retriever w/ BM25 [34] 58.31 48.39 50.73
w/ SentenceBert [33] 62.74 50.14 51.80
w/ E5 [46]) 68.42 52.84 54.31
w/ Openai-Embedding-Model 72.35 57.11 59.78
Path-RAG w/ BM25 [34] 70.34 56.11 58.77
w/ SentenceBert [33] 73.45 58.41 60.45
w/ E5 [46]) 77.93 62.74 65.23
w/ Openai-Embedding-Model 79.32 63.12 67.34
Comparison using different searching widths and depths. To explore the influence of the search width and depth on FiDeLiS’s performance, we conduct experiments under settings with depths ranging from 1 to 4 and widths from 1 to 4. As shown in Figure 2, FiDeLiS’s performance improves with the search depth and width and when the searching depth is larger than 3, FiDeLiS’s performance start to decrease on both WebQSP and CWQ. This is because only a small part of questions in these two datasets have reasoning depths greater than 3, and also FiDeLiS implements stricter rules for terminating searches using deductive reasoning. However, for searching width, we find that the performance is consistent and implies the FiDeLiS’s performance could potentially be improved with the increment of the exploration breadth. Considering the computational cost (which increases linearly with the depth), we set both the default beam width and depth as 4 without specific annotations in all the experiments.
Figure 2: Performance of FiDeLiS with different beam search widths and depths.
Comparison between Path-RAG and Vanilla Retriever. To verify the validity of the Path-RAG compared to vanilla retrievers. We conduct the experiments as shown in Figure 3. For the vanilla retriever, we concatenate each entity with its corresponding relation to form a reasoning step. We then calculated the cosine similarity between the embeddings of each reasoning step and the query embeddings to select candidate paths.
Figure 3: Comparison between Path-RAG and Vanilla Retriever.
We calculate the coverage ratio (CR) of the retrieved reasoning paths and the ground-truth reasoning paths as follows: $\text{CR}=\left(\frac{N_{\text{retrieved}}\cap N_{\text{ground-truth}}}{N_{\text{ground-truth}}}\right)\times 100\%$, where $N_{\text{retrieved}}$ is the number of retrieved reasoning paths candidates, and $N_{\text{ground-truth}}$ is the total number of ground-truth reasoning paths. We observe that compared to the vanilla retriever, Path-RAG achieves a higher CR value and shows better alignment with the ground-truth paths. It demonstrates that our method can better leverage the connections that are often missed by simpler retrieval models. This feature is critical in subsequent LLM to focus on the information of interest and arrive at a correct and reasonable answer.
Table 4: Average depths of the returned reasoning paths. The average depth of ground-truth reasoning paths of WebQSP is 2.3, for CWQ is 3.2, and for CR-LT is 4.7.
Method WebQSP CWQ CR-LT
ToG 3.1 4.1 5.2
FiDeLiS 2.4 2.8 4.6
Effectiveness of deductive-verification. To verify the effectiveness of deductive-verification, we calculate the average depths of the grounded reasoning path from ToG and our methods in Table 4. It shows that FiDeLiS consistently shows shorter and closer reasoning depths to ground-truth across all datasets compared to ToG. This implies that our method may offer more precise termination signals and potentially more accurate reasoning paths.
Error Analysis regarding Whole Path Generation.
Figure 4: Validation of whole path generation from RoG [29] over WebQSP.
To verify the faithfulness of our stepwise method, we conduct the error analysis regarding the whole reasoning path generation. Specifically, we conduct an analysis of the validity of the whole reasoning path generation using the baseline methods RoG [29]. The result is shown in Figure 4. The definition of the metric validity ratio (VR) is the ratio of reasoning steps that existed in the knowledge graph to the total number of the reasoning steps in the output reasoning path: $\text{VR}=\left(\frac{N_{\text{valid-steps}}}{N_{\text{all-steps}}}\right)\times 100\%$, where $N_{\text{valid-steps}}$ is the number of reasoning steps that existed in the knowledge graph, and $N_{\text{all-steps}}$ is the number of all the reasoning steps in the reasoning path. We note that 33% of the generated reasoning paths are invalid, 18% exhibit formatting errors, and 15% include relations that do not exist in the knowledge graph. This illustrates the limitations of embedding knowledge graphs into base LLM models, particularly in controlling the generation process.
Case Study. To verify whether FiDeLiS can conduct faithful reasoning on real cases, we also present a case study in Table 5. We can find that FiDeLiS illustrates better performance in identifying multiple possible governmental forms—Theocracy, Unitary state, and Islamic republic established in Iran in 1979, as compared to existing models such as GPT + CoT, RoG, and ToG. This observation highlights the importance of recognizing the complexity of Iran’s government system, which combines elements of religion and democracy. Different models handle this complexity with varying levels of detail. For example, GPT + CoT and RoG mainly identified Iran as an ‘Islamic Republic’, a change that occurred after the 1979 Iranian Revolution, but they missed other possible descriptions. The ToG model even gives mixed results, such as simply naming it ‘Iran’, which shows confusion in its analysis. However, our method shows a better understanding of the context and provides a more detailed and accurate set of answers. This illustrates FiDeLiS strength and efficiency in dealing with intricate questions.
Table 5: Case study of FiDeLiS. The wrong answers are highlighted with red color, and the correct answers are highlighted with blue color. In this case, FiDeLiS scored an F1 of 0.857, outperforming the RoG method (0.4).
Question What form of government is in the country that uses the Iranian Rail and was established in 1979?
Answer ’Parliamentary system’, ’Presidential system’, ’Islamic republic’, ’Unitary state’, ’Theocracy’
GPT + CoT Think step by step. First, identify the country: the country that uses the Iranian Rail is Iran. Then, retrieve the historical Context of 1979: In 1979, Iran underwent a significant transformation with the Iranian Revolution, which led to the overthrow of the Pahlavi monarchy. Third, following the revolution, the country adopted a new form of government based on Islamic law. The result was the establishment of the Islamic Republic of Iran, which combines elements of a theocracy and a democracy, where religious leaders have significant influence over the state governance. So, the government form is the Islamic Republic of Iran, established in 1979 following the Iranian Revolution.
RoG location.country.form_of_government, location.location.containedby location.country.form_of_government’ location.location.containedby $\to$ government.form_of_government.countries The results are Islamic republic.
ToG Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy $\to$ government.form_of_government.countries $\to$ Iran Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state $\to$ government.form_of_government.countries $\to$ Iran The results are Iran.
FiDeLiS Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Islamic republic Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Theocracy Iranian rial $\to$ finance.currency.countries_used $\to$ Iran $\to$ location.country.form_of_government $\to$ Unitary state The results are Theocracy, Unitary state, Islamic republic.
4 Related Work
KG-enhanced LLM. KGs have advantages in dynamic, explicit, and structured knowledge representation and techniques combining LLMs with KGs have been studied [32]. Early studies [29, 55] embed structured knowledge from KGs into the underlying neural networks during the pretraining or fine-tuning process. However, the reasoning steps generated by LMs are observed to be prone to errors, which could be non-existent in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Also, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17]. Another solutions are to keep the reasoning model as an agent to explore the external structure knowledge source [38, 9, 20, 44] ToG [38] directly employs LLMs to output scores for candidate selections from the KG. This model operates at the relation-entity level of the KG, aiming to identify relevant triples that aid the LLM in making accurate and responsible final answer predictions.
Knowledge Graph-based Question Answering. Conventional embedding-based methods represent the entities and relations in embedding space and design special model architectures (e.g., Key-Value memory networks, sequential models, and graph neural networks) to reason answers [30, 12, 52]. To integrate LLMs for KGQA, retrieval-augmented methods aim to retrieve the relative facts from the KGs to improve the reasoning performance [25, 22]. Recently, UniKGQA [21] which unifies the graph retrieval and reasoning process into a single model with LLMs, achieves STOA performance. Semantic parsing methods convert the question into a structural query (e.g., SPARQL) by LLMs, which can be executed by a query engine to reason the answers on KGs [39, 24]. However, these methods heavily rely on the quality of generated queries. If the query is not executable, no answers will be generated. DECAF [55] combines semantic parsing and LLMs reasoning to jointly generate answers, which also reach salient performance.
5 Conclusion
This paper proposes a retrieval-exploration interactive method specifically designed to enhance intermediate steps of LLM reasoning grounded by KGs. The Path-RAG module and the use of deductive reasoning as a calibration tool effectively guide the reasoning process, leading to more accurate knowledge retrieval and prevention of misleading reasoning chains. Extensive experiments demonstrate that our method, being training-free, not only reduces computational costs but also offers superior generality, consistently outperforming established strong baselines across three distinct benchmarks. We believe this study will significantly benefit the integration of LLMs and KGs, or serve as an auxiliary tool to enhance the interpretability and factual reliability of LLM outputs.
Appendix A Defintions
A.1 Defining a Valid Reasoning Trace/Path
Definition 1. A reasoning step is a pair $(r,e)$, where $r$ is the relation and $e$ is the corresponding entity.
Definition 2. A reasoning path $\mathcal{P}$ is a pair $(s,\mathcal{T})$, where $s$ is the starting entity for the reasoning path, and $\mathcal{T}$ is a sequence of reasoning steps $\mathcal{T}=\left\{t_{0},t_{1},\ldots,t_{n}\right\}$ and $t_{k}=(r_{k},e_{k})$. The $t_{k}\in\mathcal{R}$ denotes the $k$-th reasoning step in the path and $n$ denotes the length of the path.
[…]
A.2 Knowledge Graph-based Question Answering (KGQA)
In this work, we mainly focus on the question answering task based on the knowledge graph.
Knowledge Graphs (KGs) contain abundant factual knowledge in the form of a set of triples: $\mathcal{G}=$$\left\{\left(e,r,e^{\prime}\right)\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\right\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations, respectively.
Knowledge Graph Question Answering (KGQA) is a typical reasoning task based on KGs. Given a natural language question $q$ and a $\mathrm{KG}$ $\mathcal{G}$, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ based on knowledge from $\mathcal{G}$, i.e., $a=f(q,\mathcal{G})$. Following previous works [29, 28], we assume the entities $e_{q}\in\mathcal{T}_{q}$ mentioned in $q$ and answers $a\in\mathcal{A}_{q}$ are labeled and linked to the corresponding entities in $\mathcal{G}$, i.e., $\mathcal{T}_{q},\mathcal{A}_{q}\subseteq\mathcal{E}$.
Appendix B Prompt List
In this section, we show all the prompts that need to be used in the main experiments. The In-Context Few-shot refers to the few-shot examples we used for in-context learning.
B.1 Plan-and-solve
You are a helpful assistant designed to output JSON that aids in navigating a knowledge graph to answer a provided question. The response should include the following keys:
(1) ’keywords’: an exhaustive list of keywords or relation names that you would use to find the reasoning path from the knowledge graph to answer the question. Aim for maximum coverage to ensure no potential reasoning paths will be overlooked;
(2) ’planning_steps’: a list of detailed steps required to trace the reasoning path with. Each step should be a string instead of a dict.
(3) ’declarative_statement’: a string of declarative statement that can be transformed from the given query, For example, convert the question ’What do Jamaican people speak?’ into the statement ’Jamaican people speak *placeholder*.’ leave the *placeholder* unchanged; Ensure the JSON object clearly separates these components.
In-Context Few-shot
Q: {Query}
A:
B.2 Deductive-verification
You are asked to verify whether the reasoning step follows deductively from the question and the current reasoning path in a deductive manner. If yes return yes, if no, return no".
In-Context Few-shot
Whether the conclusion ’{declarative_statement}’ can be deduced from ’{parsed_reasoning_path}’, if yes, return yes, if no, return no.
A:
B.3 Adequacy-verification
You are asked to verify whether it’s sufficient for you to answer the question with the following reasoning path. For each reasoning path, respond with ’Yes’ if it is sufficient, and ’No’ if it is not. Your response should be either ’Yes’ or ’No’.
In-Context Few-shot
Whether the reasoning path ’{reasoning_path}’ be sufficient to answer the query ‘{Query}’, if yes, return yes, if no, return no.
A:
B.4 Beam Search
Given a question and the starting entity from a knowledge graph, you are asked to retrieve reasoning paths from the given reasoning paths that are useful for answering the question.
In-Context Few-shot
Considering the planning context {plan_context} and the given question {Query}, you are asked to choose the best {beam_width} reasoning paths from the following candidates with the highest probability to lead to a useful reasoning path for answering the question. {reasoning_paths}. Only return the index of the {beam_width} selected reasoning paths in a list.
A:
B.5 Reasoning
Given a question and the associated retrieved reasoning path from a knowledge graph, you are asked to answer the following question based on the reasoning path and your knowledge. Only return the answer to the question.
In-Context Few-shot
Question: {Query}
Reasoning path: {reasoning_path}
Only return the answer to the question.
A:
Appendix C Potential Impacts and Limitations
The proposed retrieval-exploration interactive method, FiDeLiS, holds the potential to significantly enhance the performance of LLMs by addressing the issue of hallucinations and enabling deep, responsible reasoning. By integrating KGs with LLMs, the method offers more accurate knowledge retrieval, leveraging deductive reasoning capabilities to guide the reasoning process and avoid logical fallacies. The stepwise and generalizable approach, along with the use of deductive verification as a stop criterion, may reduce unnecessary computation and misleading reasoning chains. Additionally, the training-free nature of this method combined with its lower computational cost could be easily acted as a plugin-in for other methods/frameworks.
However, there are limitations to consider. As the method is still in its early stages of development, there may be unforeseen challenges or drawbacks when applied to real-world scenarios. The reliance on external knowledge graphs means that the quality and coverage of these resources can impact the overall effectiveness of the method. There may also be computational constraints when working with very large or complex graphs. Additionally, while the method has shown promise in benchmarks, its performance in more diverse or specialized tasks is yet to be evaluated. Despite these challenges, the research presents a promising direction for improving the reasoning capabilities of LLMs.
Appendix D Efficiency of FiDeLiS
The bottleneck of computation is the beam search process, which contributes to $N*D$ times LLM calling, where $D$ is the depth (or equivalently length) of the reasoning path, and $N$ is the width of the beam-search (how many paths are remained in the pool in each iteration). Specifically, we need to call $ND+D+C$ times LLM for each sample question, where $C$ is a constant (equals to $1$ if there is no error occurs when calling the API). [38] illustrates that the computational efficiency can be alleviated by replacing LLMs with small models such as BM25 and Sentence-BERT for the beam search decision since the small models are much faster than LLM calling. In this way, we can reduce the number of LLM calling from $ND+D+C$ to $D+C$. However, [38] shows that this optimization sacrifices the accuracy due to the weaker scoring model in decision making.
Table 6: Average number of LLM calls per question.
Dataset WebQSP CWQ CR-LT
Average LLM Calls 10.7 15.2 17.3
We noted that $ND+D+C$ is the maximal computational complexity. In most cases, FiDeLiS does not need $ND+D+C$ LLM calls for a question because the whole reasoning process might be early stopped before the maximum reasoning depth $D$ is reached if LLM determines the query can be deductive reasoning by the current retrieved reasoning paths. As an illustration, Table 6 shows the average numbers of LLM calls per question needed by FiDeLiS on different datasets. It can be seen that in three KGQA datasets, the average numbers of LLM calls (ranging from ) are smaller than $21$, which is the theoretical maximum number of LLM calls calculated from $ND+D+C$ when $N=4$ and $D=4$. We can also see that this average number gets even smaller for dataset covering a lot of single-hop reasoning questions, such as WebQSP.
Table 7: Average token usage of LLM calls per question.
Dataset WebQSP CWQ CR-LT
Average Token Usage 2,452 2,741 1,824
Cost efficiency. Another benefit of FiDeLiS is the cost efficiency. We control the token usage for each LLM calling through Path-RAG module. Specifically, Path-RAG only retrieves the top-n reasoning steps as the candidates for each beam search decision making, before making calls to the LLM. This approach ensures token expenditure without compromising the model’s performance, especially in some cases where hundreds of neighbors exist around a entity node. Table 7 shows the average token usage per question needed by FiDeLiS on different datasets. It can be seen that in three KGQA datasets, the numbers of token usages are ranging from 1,824 to 2,751, which could significantly reduce the cost for running the method.
Appendix E Variants of Path-RAG
Variant-1: To enhance the scoring function for reasoning in a knowledge graph and ensure each entity and relation are consistently linked, we can modify the scoring function to more explicitly account for the connectivity and interdependences between entities and relations throughout the path. This can be done by incorporating an accumulative sum that keeps track of previous scores and emphasizes path completeness. The revised formula introduces variables for the accumulated scores of entities and relations, which will help ensure all links in the path are considered:
$S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}\left(S_{r}^{j}+S_{e}^{j}\right)+\beta\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$ (8)
where: (1) $S_{r}^{i}$ and $S_{e}^{i}$ denote the relevance score of the current step relation and the corresponding neighboring entity, respectively. (2) $S_{r}^{j}$ and $S_{e}^{j}$ denote the relevance score of the 1-hop relation and the corresponding neighboring entity from node $N_{i}$. (3) The new term, $\sum_{k=0}^{i-1}\left(S_{r}^{k}+S_{e}^{k}\right)$, accumulates the scores from all previous steps from the start up to the step before the current, enhancing the influence of the entire path on the decision of the next step. (4) $\alpha$ and $\beta$ are weight factors that control the influence of the maximum score of neighboring nodes and the accumulated score to adjust the balance in prioritizing extended paths versus immediate gains. This adjustment helps to: firstly, reward paths that maintain strong links throughout, discouraging breaks or weakly connected steps in the reasoning process; secondly, enhance the depth of consideration in the decision-making process, looking both at immediate and long-term implications of each step.
Variant-2: We also consider to incorporate the randomness into the scoring function to balance the quality-diversity trade-off in searching for better reasoning paths as follows:
$S(p)=\lambda_{1}\left(S_{r}^{i}+S_{e}^{i}+\alpha\sum_{j\in N_{i}}(S_{r}^{j}+S_{e}^{j})\right)+\lambda_{2}\sum_{k=0}^{i-1}\left(D(p,P_{k})\right)$ (9)
where $D(p,P_{\text{sel}})$ refers to the distance between the selected path with the grounded reasoning paths. while $\lambda_{1}$ and $\lambda_{2}$ are weighting factors that balance the importance of internal quality metrics against diversity metrics.
Appendix F Additional Related Works: Reasoning with LLM Prompting
With the development of LLMs, many creative ways to leverage LLMs are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation.
End-to-End methods [4, 15] aims to leverage LLMs to generate final answers directly, often done by providing a task description and/or a few examples for in-context learning. While it offers convenience, however, it suffers from the un-interpretability of the generation, and the lack of explicit steps and reliance solely on the LLMs’ training data may result in un-robustness and exhibit sensitivity to slight input variations.
Chain-of-Thought methods [48, 23, 8] emphasize breaking down complex reasoning into a series of intermediate steps to support complex reasoning. However, CoT suffers from unreliability and uncontrollability since the generated reasoning steps may not always align with the intended logical progression or may produce incorrect or inconsistent answers when multiple questions are posed consecutively.
Plan-and-solve methods [45] prompts LLMs to generate a plan and conduct reasoning based on it. DecomP [12] prompts LLMs to decompose the reasoning task into a series of sub-tasks and solve them step by step.
Semantic Parsing/Code Generation methods [5, 53, 7] leverages LLMs to convert natural language queries into executable code or structured representations. This approach enables more precise control over the output and facilitates better interpretability. However, it has its limitations, such as the limited coverage of programming language grammar and semantics. This restricts the model’s ability to handle complex programming tasks and may require additional techniques to handle out-of-domain queries effectively. [5] propose to leverage LLMs to bridge the out-of-domain queries, however, the form of knowledge injection still lacks a thorough exploration and the trigger functions are still naive for complex queries. However, the problem of hallucinations and lack of knowledge affect the faithfulness of LLMs’ reasoning. ReACT [51] treats LLMs as agents, which interact with the environment to get the latest knowledge for reasoning. To explore faithful reasoning, FAME [16] introduces the Monte-Carlo planning to generate faithful reasoning steps. RR [13] and KD-CoT [44] further retrieve relevant knowledge from KGs to produce faithful reasoning plans for LLMs.
Appendix G Experiment Details
G.1 Baselines
RoG [29]: It aims to embed structure knowledge from KGs into the underling neural networks during the pretraining and fine-tuning process. However, the controllablity of the reasoning paths generation is hard to achieve since the decoding process can be shortsighted, i.e., it is difficult to tell whether the current operation/decision would lead to a malicious operation several steps later. Also, the reasoning steps generated by LLMs are observed to be prone to errors, which could be non-exist in the knowledge graph, and can subsequently lead to incorrect reasoning during inference. Additionally, KG embedded in LLM sacrifices its own nature of explainability in knowledge reasoning and efficiency in knowledge updating [17]. Compared to RoG, FiDeLiS propose to employ beam search tailored for intermediate reasoning steps and guide the beam searching process by controlling the error of each reasoning step to prevent accumulation thought the chaining.
ToG [38]: It aims step-wise grounding to elicit more accurate results. However, it directly employs LLMs to output scores for candidate selections from the KG, which require an LLM/LM to iteratively score candidate plans, which are resource-consuming in terms of both time and computing. Also, the performance of the proposed ToG [38] seems rely heavily on the use of very strong LLMs (including Codex [3], GPT-4 [31]) to do the final reasoning; and the true potential of enabling few-shot grounded reasoning path understanding has yet to be fully realized.
G.2 Datasets
Table 8: Statistics of the number of answers for questions in WebQSP and CWQ.
Dataset #Ans = 1 2 $\geq$ #Ans $\leq$ 4 5 $\geq$ #Ans $\leq$ 9 #Ans $\geq$ 10
WebQSP 51.2% 27.4% 8.3% 12.1%
CWQ 70.6% 19.4% 6% 4%
Table 9: Statistics of the question hops in WebQSP, CWQ and CR-LT-KGQA.
Dataset 1 hop 2 hop $\geq$ 3 hop
WebQSP 65.49 % 34.51% 0.00%
CWQ 40.91 % 38.34% 20.75%
CR-LT 5.31 % 43.22% 51.57%
Figure 5: Distribution of CR-LT-KGQA dataset.
Motivation of CR-LT-KGQA The motivation for evaluating over CR-LT-KGQA is that the majority of existing KGQA datasets, including WebQSP and CWQ, predominantly feature popular entities. These entities are well-represented in the training corpora of LLMs, allowing to often generate correct answers based on their internal knowledge, potentially without external KGs. Moreover, since WebQSP and CWQ have been available for several years, there is a likelihood that their data might have been incorporated into the pre-training corpora of LLMs, further reducing the need for external KGs during question-answering. Against this backdrop, we specifically chose CR-LT-KGQA for our evaluation as it includes queries related to obscure or long-tail entities, the distribution of the frequency and entity popularity of CR-LT is shown as Figure 5. Such scenarios are where KGs play a critical role because they provide a reliable source of verifiable information, crucial when LLMs encounter entities not well covered in their training data. By testing our methods on CR-LT-KGQA, we aim to explore how effective LLMs can operate when combined with KGs, especially in contexts involving less common knowledge domains where LLM performance typically declines. This evaluation helps us understand the extent to which KGs remain necessary for supporting LLMs in a diverse range of query scenarios.
Table 10: Comparison of FiDeLiS using different backbone LLMs.
Backend Models Methods WebQSP CWQ CR-LT
Hits@1 (%) F1 (%) Hits@1 (%) F1 (%) Acc (%)
Llama-2-13B FiDeLiS 72.34 69.78 58.41 54.78 60.87
Mistral-7B FiDeLiS 74.11 70.23 60.71 56.87 63.12
Example of Deductive Verification Question: “Who is the ex-wife of Justin Bieber’s father?" After one round of beam searching, the current reasoning path is “Justin_bieber $\to$ people.person.father $\to$ Jeremy_bieber", and the next step candidates are “1. people.married_to.person $\to$ Erin Wagner"; 2. people.person.place_of_birth $\to$ US, $\ldots$". The deductive reasoning can be formulated as: Premises: Justin_bieber $\to$ people.person.father —> Jeremy_bieber (from the current reasoning path) Jeremy_bieber $\to$ people.married_to.person $\to$ Erin Wagner (from the next step candidates) Conclusion: Erin Wagner is the ex-wife of Justin Bieber’s father. (Using LLM zero-shot to reformat the question to a cloze filling, and use the last entity from the next step candidate"Erin Wagner” to fill the cloze) The prompt will be whether the conclusion can be deduced from the given premises, if yes, return “yes”, it no, return “no".
Appendix H Algorithm Description
Algorithm 1 Path-RAG Initialization and Retrieval Process
1:Initialization:
2:for all $e_{i}\in\mathcal{E},r_{i}\in\mathcal{R}$ do
3:     $z_{e}^{i}=\operatorname{LM}(e_{i})$triangleright Embed entities
4:     $z_{r}^{i}=\operatorname{LM}(r_{i})$triangleright Embed relations
5:end for
6:Populate nearest neighbor index with $\{z_{e}^{i}\}$ and $\{z_{r}^{i}\}$triangleright Facilitate retrieval
7:procedure Retrieve(query $q$)
8:     $\mathcal{K}_{i}=\operatorname{LM}(\text{`prompt'},q)$triangleright Generate keywords
9:     for all $k_{i}^{m}\in\mathcal{K}_{i}$ do
10:         $k_{i}\leftarrow\text{concatenate}(k_{i}^{m})$
11:         $z_{k}=\operatorname{LM}(k_{i})$triangleright Embed concatenated keywords
12:         $\mathcal{E}_{k}=\operatorname{argtopk}_{i\in\mathcal{E}}\cos(z_{k},z_{e}^{i})$triangleright Retrieve top-k entities
13:         $\mathcal{R}_{k}=\operatorname{argtopk}_{i\in\mathcal{R}}\cos(z_{k},z_{r}^{i})$triangleright Retrieve top-k relations
14:     end for
15:     return $\mathcal{E}_{k},\mathcal{R}_{k}$
16:end procedure
17:procedure ScorePath($\mathcal{E}_{k},\mathcal{R}_{k}$)
18:     Initialize $\text{Score}\leftarrow 0$
19:     for each $e_{k}\in\mathcal{E}_{k}$ and $r_{k}\in\mathcal{R}_{k}$ do
20:         Calculate $S_{e}^{i},S_{r}^{i}\leftarrow\cos(z_{k},z_{e}^{i}),\cos(z_{k},z_{r}^{i})$triangleright Compute similarity scores
21:         $S(p)=S_{r}^{i}+S_{e}^{i}+\alpha\max_{\forall j\in N_{i}}(S_{r}^{j}+S_{e}^{j})$triangleright Score path using Eq. 4
22:         $\text{Score}\leftarrow\max(\text{Score},S(p))$triangleright Update max score
23:     end for
24:     return $\text{Score},p$
25:end procedure
Algorithm 2 Deductive-Verification Guided Beam Search
1:User query $x$, Beam width $B$
2:Reasoning path $s^{1:T}$
3:Initialize $\mathcal{H}_{0}=\{\emptyset\}$
4:Utilize LLM to generate from $x$:
5:      Planning steps.
6:      Declarative statement $x^{\prime}$.
7:for $t=1$ to $T$ do
8:     for each $h\in\mathcal{H}_{t-1}$ do
9:         Generate possible next steps $s^{t}\in\mathcal{S}$ using Path-RAG.
10:         for each $s^{t}$ do
11:              Compute $C(x^{\prime},s^{t},s^{1:t-1})$ using LLM:
12:                    $C(x^{\prime},s^{t},s^{1:t-1})=\begin{cases}1&\text{if }x^{\prime}\text{ can bededuced from }s^{t}\text{ and }s^{1:t-1},\\ 0&\text{otherwise.}\end{cases}$
13:              if $C(x^{\prime},s^{t},s^{1:t-1})=1$ then
14:                  Append $s^{t}$ to $h$ to form new hypothesis $h^{\prime}$.
15:                  Add $h^{\prime}$ to $\mathcal{H}_{t}$.
16:              end if
17:         end for
18:     end for
19:     $\mathcal{H}_{t}=\text{Top}_{B}(\mathcal{H}_{t})$ based on scoring function (like plausibility or likelihood).
20:end for
21:return the best hypothesis from $\mathcal{H}_{T}$.